OpenStudy (anonymous):

7/x-1 + 4/x+1=3x/x^2-1 7/x-1 + 4/x+1=3x/x^2-1 @Mathematics

OpenStudy (agreene):

\[\frac{7}{x-1}+\frac{4}{x+1}=\frac{3x}{x^2-1}\] is that the question?

OpenStudy (anonymous):

yes. I have no clue how to type it in the form like that lol

OpenStudy (anonymous):

i know the lcd is x^2-1

OpenStudy (anonymous):

ive gotten to where i multiply that on both sides

OpenStudy (agreene):

hehe that's fine... its just hard to tell what things are with no parenthesis. to solve: I would start by factoring the denominator on the right and getting: \[\frac{7}{x-1}+\frac{4}{x+1}=\frac{3x}{(x-1)(x+1)}\] So, now we can multiply both sides by that denominator and we get: \[7(x+1)+4(x-1)=3x\] now we can distribute the 7 and 4: \[7x+7+4x-4=3x\] combine likes: \[11x+3=3x\] subtract 3x and 3 from both sides \[8x=-3\] divide by 8 \[x=-\frac{3}{8}\]

OpenStudy (anonymous):

i got it right! thanks! i get lost when it comes to multiplying

OpenStudy (anonymous):

i have like 5 questions im having to make sure i have correct before i turn my work in lol